HETAV PATEL

Grade 11,

Solve the integral

Solve the integral


 


4063-1727_3899_indefinite1.bmp

Grade:12

1 Answers

Ramesh V
14 years ago

you can write separately as:

= x/(x+1)2 - 3/(x+1)2  + 2/x.(x+1)2

= (x+1-1)/(x+1)2 - 3/(x+1)2  + 2/x.(x+1)2

= 1/(x+1) - 4/(x+1)2  + 2/x.(x+1)2

= 1/(x+1) - 4/(x+1)2  + 2/x - 2/(x+1) - 2/(x+1)2

= 2/x - 1/(x+1) - 6/(x+1)2

on integrating gives

2 lnx - ln(x+1) +6/(x+1) +C